### v1

@@ Appendix A Properties of Gaussian matrices
$\mathbb{P}\left\{\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}>\sqrt{\frac{3\operatorname{tr}(\widetilde{\bm{K}}_{11}^{-1})}{p+1}}t\right\}$ $\leq t^{-p},$ (36a)
$\mathbb{P}\left\{\|\bm{\Omega}_{1}^{\dagger}\|_{2}>\frac{e\sqrt{(k+p)\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}}}{p+1}t\right\}$ $\leq t^{-(p+1)},$ (36b)
$\mathbb{P}\left\{\|\bm{\Omega}_{1}^{\dagger}\|_{(4)}>\frac{e\sqrt{(k+p)\|\widetilde{\bm{K}}_{11}^{-1}\|_{\textup{F}}}}{p+1}t\right\}$ $\leq t^{-(p+1)}.$ (36c)
Proof A.8.
Note that Eq. 36a is a restatement of [8, Lem. 3]. Because $\bm{\Omega}_{1}=\widetilde{\bm{K}}_{11}^{1/2}\bm{\Psi}$, where $\bm{\Psi}$ is a standard Gaussian matrix, it follows that $\|\bm{\Psi}^{\dagger}\widetilde{\bm{K}}_{11}^{-1/2}\|_{(s)}\leq\|\bm{\Psi}^{\dagger}\|_{2}\|\widetilde{\bm{K}}_{11}^{-1/2}\|_{(s)}$ for any Schatten-$s$ norm. Moreover, the combination with [20, Prop. 10.4] implies
@@
which yield the bounds Eqs. 36b and 36c using $\|\widetilde{\bm{K}}_{11}^{-1/2}\|^{2}_{2}=\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}$ and $\|\widetilde{\bm{K}}_{11}^{-1/2}\|^{2}_{(4)}=\|\widetilde{\bm{K}}_{11}^{-1}\|_{\textup{F}}$, respectively.

### v2

@@ Appendix A Properties of Gaussian matrices
(40a) $\mathbb{P}\left\{\|\bm{\Omega}_{1}^{\dagger}\|_{\textup{F}}>\sqrt{\frac{3\operatorname{tr}(\widetilde{\bm{K}}_{11}^{-1})}{p+1}}t\right\}$ $\leq t^{-p},$
(40b) $\mathbb{P}\left\{\|\bm{\Omega}_{1}^{\dagger}\|_{2}>\frac{e\sqrt{(k+p)\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}}}{p+1}t\right\}$ $\leq t^{-(p+1)},$
(40c) $\mathbb{P}\left\{\|\bm{\Omega}_{1}^{\dagger}\|_{(4)}>\frac{e\sqrt{(k+p)\|\widetilde{\bm{K}}_{11}^{-1}\|_{\textup{F}}}}{p+1}t\right\}$ $\leq t^{-(p+1)}.$
Proof A.6.
Note that Eq. 40a is a restatement of [8, Lem. 3]. Because $\bm{\Omega}_{1}=\widetilde{\bm{K}}_{11}^{1/2}\bm{\Psi}$, where $\bm{\Psi}$ is a standard Gaussian matrix, it follows that $\|\bm{\Psi}^{\dagger}\widetilde{\bm{K}}_{11}^{-1/2}\|_{(s)}\leq\|\bm{\Psi}^{\dagger}\|_{2}\|\widetilde{\bm{K}}_{11}^{-1/2}\|_{(s)}$ for any Schatten-$s$ norm. Moreover, the combination with [21, Prop. 10.4] implies
@@
which yield the bounds Eqs. 40b and 40c using $\|\widetilde{\bm{K}}_{11}^{-1/2}\|^{2}_{2}=\|\widetilde{\bm{K}}_{11}^{-1}\|_{2}$ and $\|\widetilde{\bm{K}}_{11}^{-1/2}\|^{2}_{(4)}=\|\widetilde{\bm{K}}_{11}^{-1}\|_{\textup{F}}$, respectively.